A lock consists of 3 dials, where each dial has 4 letters. What is the probability of guessing the right combination in one try?

1. 1/7

2. 1/12

3. 1/24

4. 1/64
Answer #1

Solution:

A lock consists of 3 dials, where each dial has 4 letters.

Total number of possible combinations = 3 * 4 = 12

Sample space contains 12 elements.

For a lock , only one combination is correct.

So ,

probability of guessing the right combination in one try = 1/12

Option 2. 1/12